OpenStudy (anonymous):

Jill made a trip to the train station and back. the trip there took 2 hours and the trip back took 3 hours. she averaged 20 mph faster on the trip there than on the way back. find jills average speed on the outbound trip. please/thanks

OpenStudy (anonymous):

let her speed be x and x-20 units respectively by the problem we have 2x=3(x-20) or 2x=3x-60 or x=60 thus she went there with speed of 60 mph and returned back with 40mph

OpenStudy (anonymous):

thank you!